# Beyond the Hodge Theorem: curl and asymmetric pseudodifferential projections

### v3

@@ Abstract
We develop a new approach to the study of spectral asymmetry. Working with the operator $\operatorname{curl}:=*\mathrm{d}$ on a connected oriented closed Riemannian 3-manifold, we construct, by means of microlocal analysis, the asymmetry operator — a scalar pseudodifferential operator of order $-3$. The latter is completely determined by the Riemannian manifold and its orientation, and encodes information about spectral asymmetry. The asymmetry operator generalises and contains the classical eta invariant traditionally associated with the asymmetry of the spectrum, which can be recovered by computing its regularised operator trace. Remarkably, the whole construction is direct and explicit.
@@ 1 Main results
Let $(M,g)$ be a connected closed Riemannian manifold of dimension $d\geq 2$ and let $\Omega^{k}(M)$ be the Hilbert space of real-valued $k$-forms, $11\leq k\leq d-$ 1. Hodge’s Theorem [36, Corollary 3.4.2] tells us that $\Omega^{k}(M)$ decomposes into a direct sum of three orthogonal closed subspaces
@@
To the best of our knowledge, the only paper dedicated to the study of the spectrum of $\operatorname{curl}$ on a closed oriented Riemannian manifold is [7]. In [7] the author examined the qualitative properties of the spectrum of $\operatorname{curl}$ depending on the dimension $d$ and established one-term asymptotic formulae (Weyl law) for the (global) eigenvalue counting functions, alongside a number of explicit examples. Although not very close in spirit to the aims of the current paper, we also refer the reader to [25, 26] examining the first eigenvalue of $\operatorname{curl}$ on domains in $\mathbb{R}^{3}$.
@@
We will show in Section 2 — more precisely, in Theorem 2.1 — that the differential expression (1.5) gives rise to a self-adjoint operator (denoted by the same symbol) in the space of coexact 1-forms $\delta\Omega^{2}(M)$. The latter has discrete spectrum, accumulating to both +$\infty$+ ∞ and $\infty$- ∞, and not necessarily symmetric about zero. As a way of illustrating that spectral asymmetry actually occurs, in Appendices B and C we write down explicitly the spectrum of $\operatorname{curl}$ on a Berger sphere and compute the corresponding (classical) eta invariant.
Note that the definition of $\operatorname{curl}$ does not involve the concept of connection. In this respect, $\operatorname{curl}$ is one of the most fundamental operators of mathematical physics, like the Laplace–Beltrami operator. Furthermore, $\operatorname{curl}$ lies at the heart of (homogeneous) Maxwell’s equations. Namely, seeking solutions harmonic in time reduces Maxwell’s equations on $M\times\mathbb{R}$ to the spectral problem for $\operatorname{curl}$. And the solution of the Cauchy problem for Maxwell’s equations can be expressed in terms of eigenvalues and eigenforms of $\operatorname{curl}$, see formula (D.9). Finally, it is worth emphasising that the sign of the eigenvalues of curl has a physical meaning: it coincides with the sign of electromagnetic chirality of time-harmonic polarised solutions of Maxwell’s equations generated by the corresponding $\operatorname{curl}$ eigenpair (see (D.6), (D.7)). We refer the reader to Appendix D for further details and a self-contained exposition of this fact. These are but few of the reasons why a more detailed examination of the spectrum of curl — and in particular of its spectral asymmetry — is very timely, if not overdue.
@@
[…]
@@
    The operators $P_{0}$, $P_{+}$ and $P_{-}$ are pseudodifferential of order zero.
@@
    $[(P_{0})_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)=\|\xi\|^{-2}\,\xi_{\alpha}\,g^{\beta\gamma}(x)\,\xi_{\gamma}\,,$ (1.8)
    $[(P_{\pm})_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)=\frac{1}{2}\left[\delta_{\alpha}{}^{\beta}-[(P_{0})_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)\pm i\,\|\xi\|^{-1}\,E_{\alpha}{}^{\gamma\beta}(x)\,\xi_{\gamma}\right],$ (1.9)
@@
Remarkably, the pseudodifferential projections $P_{0}$ and $P_{\pm}$ can be constructed explicitly. In Proposition 5.3 we will provide an algorithm for the calculation of their full symbols. The latter proposition and Theorem 1.3 give a detailed description of the structure of the projection operators $P_{0}$ and $P_{\pm}$, which is our first main result.
@@
#{positive eigenvalues}#{negative $\#\{\text{positive eigenvalues}\}\,-\,\#\{\text{negative eigenvalues}\}=\operatorname{Tr}(P_{+}-P_{-}).$ (1.12)
@@
The operator $A$ is prima facie a pseudodifferential operator of order 0. However, taking the matrix trace brings about unexpected cancellations, so that the operator $A$ turns out, in fact, to be of order $-3$. This is our second main result.
@@
The asymmetry operator $A$ is a pseudodifferential operator of order $-3$ and its principal symbol reads
@@
A sufficient condition for a self-adjoint pseudodifferential operator to be of trace class and to have continuous integral kernel is for its order to be strictly less than $-d$, where $d$ is the dimension of the manifold. See also [49, §12.1]. In the case at hand, the operator $A$ has order $-3$ and the dimension of the manifold is 3, so we are looking at a borderline situation. One would hope that with a mild regularisation one could cross the finish line and be able to define a notion of regularised trace.
@@
Pseudodifferential projections are very natural tools in the study of the spectral properties of systems. At first glance, one may be tempted to think they are unnecessarily complicated, especially for operators acting on trivial bundles, in that one could try and diagonalise the operator along the lines of [13] instead, thus reducing the problem at hand to the examination a collection of scalar operators. However, after examining the matter more carefully, diagonalisation turns out to be topologically obstructed in numerous examples of physical relevance. In particular, it was shown in [15, Section 3.3] that for the case of the operator $\operatorname{curl}$ one cannot even diagonalise the principal symbol $\operatorname{curl}_{\mathrm{prin}}$ (2.1) in the whole cotangent fibre at a given point of $M$, let alone the operator itself.
@@
In Section 2 we put $\operatorname{curl}$ on a rigorous operator-theoretic footing, and state and prove its basic properties. In doing so, we rely on the auxiliary operator $\operatorname{curl}_{E}$, called extended curl, which can be viewed as an elliptic “extension" of $\operatorname{curl}$. Although most of the results from this section are known at least in some form, detailed proofs, often hard to find in the literature, are provided here in full and in a self-contained fashion, for the convenience of the reader and future reference.
@@
In Section 5 we study the pseudodifferential projections $P_{0}$, $P_{+}$ and $P_{-}$, and provide an explicit algorithm for the construction of their full symbols. This section contains the proof of Theorem 1.3.
In Section 6 we define the asymmetry operator $A$ (subsection 6.1), prove that it is a pseudodifferential operator of order $-3$ (subsection 6.2), and compute its principal symbol $A_{\mathrm{prin}}$ (subsection 6.3). This section establishes Theorem 1.4.
@@
The paper is complemented by a list of notation given at the end of this section, and by six appendices. In Appendix A we summarise our notation and conventions on exterior calculus. Appendices B and C are concerned with spectral theory on Berger 3-spheres: in Appendix B we recall the definition of a Berger sphere and list the eigenvalues of the Laplace–Beltrami operator, whereas in Appendix C we list the eigenvalues of $\operatorname{curl}$ on a Berger sphere as an illustration of spectral asymmetry, and compute the eta invariant by explicitly examining the eta function for curl. In Appendix D we examine the relation between the spectral problem for curl and solutions of Maxwell’s equations, and discuss the physical meaning of the sign of the eigenvalues of $\operatorname{curl}$. In Appendix E we derive series expansions for the parallel transport maps. Finally, in Appendix F we independently verify formula (1.13).
@@
[…]
$A$ Asymmetry operator, Definition 6.1
@@
$\mathfrak{a}(x,y)$ Integral kernel of the asymmetry operator $A$
@@
$\operatorname{curl}_{E}$ Extended curl, Definition 2.3
$\operatorname{curl}_{E,\mathrm{d}},\operatorname{curl}_{E,\delta},\operatorname{curl}_{E,\mathcal{H}}$ Orthogonal summands of $\operatorname{curl}_{E}$, Lemma 2.6
@@
$e_{j}{}^{\alpha}(x)$, $e^{k}{}_{\beta}(x)$ Framing and dual framing (3.8)
$\widetilde{e}_{j}{}^{\alpha}(x)$, $\widetilde{e}^{k}{}_{\beta}(x)$ Levi-Civita framing and dual Levi-Civita framing, Definition 3.5
@@
$(\mu_{j},f_{j})$, $j=0,1,2,\ldots$ Eigensystem for $-\Delta$
@@
$Q_{\mathrm{prin},s}$ Principal symbol of $Q$, a pseudodifferential operator of order $-s$
@@
$TM$, $T^{*}M$ Tangent and cotangent bundle
@@
$Z$ Parallel transport map (4.6), see also Appendix E
@@ 2 The operator curl
In this section we put $\operatorname{curl}$ on a rigorous mathematical foundation, defining it within the framework of the theory of unbounded self-adjoint operators in Hilbert spaces. This would provide the reader with an account of the fundamental properties of $\operatorname{curl}$, in a self-contained fashion.
@@
$[\operatorname{curl}_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)=-i\,E_{\alpha}{}^{\beta\gamma}(x)\,\xi_{\gamma}\,,$ (2.1)
where the tensor $E$ is defined in accordance with (1.10). Although $\xi$ is not a 1-form in its own right, at a formal level the RHS of (2.1) can be rewritten, concisely, as $-i*\xi$. A straightforward calculation shows that
@@
The eigenvalues of $(\operatorname{curl})_{\mathrm{prin}}$ are simple and read
@@ 2.1 Operator theoretic definition and main properties
The following theorem, the main result of this section, establishes the basic properties of the operator (2.4). It essentially tells us that, despite not being elliptic, $\operatorname{curl}$ enjoys numerous properties characteristic of elliptic operators.
@@
    The spectrum of $\operatorname{curl}$ is discrete and accumulates to +$\infty$+ ∞ and $\infty$- ∞.
@@
Let us point out, once again, that throughout this paper $\operatorname{curl}$ comes in two guises: as a differential expression (1.5) and as a self-adjoint operator (2.4). Both will be denoted by “$\operatorname{curl}$”; which is which will be clear from the context.

### v4

@@ Abstract
We develop a new approach to the study of spectral asymmetry. Working with the operator $\operatorname{curl}:=*\mathrm{d}$ on a connected oriented closed Riemannian 3-manifold, we construct, by means of microlocal analysis, the asymmetry operator — a scalar pseudodifferential operator of order 3-3. The latter is completely determined by the Riemannian manifold and its orientation, and encodes information about spectral asymmetry. The asymmetry operator generalises and contains the classical eta invariant traditionally associated with the asymmetry of the spectrum, which can be recovered by computing its regularised operator trace. Remarkably, the whole construction is direct and explicit.
@@ 1 Main results
Let $(M,g)$ be a connected closed Riemannian manifold of dimension $d\geq 2$ and let $\Omega^{k}(M)$ be the Hilbert space of real-valued $k$-forms, $1\leq k\leq d-1$. Hodge’s Theorem [42, Corollary 3.4.2] tells us that $\Omega^{k}(M)$ decomposes into a direct sum of three orthogonal closed subspaces
@@
To the best of our knowledge, one of the few papers dedicated to the study of the spectrum of $\operatorname{curl}$ on a closed oriented Riemannian manifold — at least, in the way it is understood in the current paper — is [7]. In [7] the author examined the qualitative properties of the spectrum of $\operatorname{curl}$ depending on the dimension $d$ and established one-term asymptotic formulae (Weyl law) for the (global) eigenvalue counting functions, alongside a number of explicit examples. Although not entirely aligned in spirit with the aims of the current paper, we refer the reader to [21, 41, 27, 29, 28, 33, 43, 47, 48] and references therein, examining the spectrum of $\operatorname{curl}$ (or $\operatorname{curl}^{2}$) from different perspectives.
@@
We will show in Section 2 — more precisely, in Theorem 2.1 — that the differential expression (1.5) gives rise to a self-adjoint operator (denoted by the same symbol) in the space of coexact 1-forms $\delta\Omega^{2}(M)$. The latter has discrete spectrum, accumulating to both $\infty$ and -$\infty$, and not necessarily symmetric about zero. As a way of illustrating that spectral asymmetry actually occurs, in Appendices C and D we write down explicitly the spectrum of $\operatorname{curl}$ on a Berger sphere and compute the corresponding (classical) eta invariant.
Note that the definition of $\operatorname{curl}$ does not involve the concept of connection. In this respect, $\operatorname{curl}$ is one of the most fundamental operators of mathematical physics, like the Laplace–Beltrami operator. Furthermore, $\operatorname{curl}$ lies at the heart of (homogeneous) Maxwell’s equations. Namely, seeking solutions harmonic in time reduces Maxwell’s equations on $M\times\mathbb{R}$ to the spectral problem for $\operatorname{curl}$. And the solution of the Cauchy problem for Maxwell’s equations can be expressed in terms of eigenvalues and eigenforms of $\operatorname{curl}$, see formula (E). Finally, it is worth emphasising that the sign of the eigenvalues of curl has a physical meaning: it coincides with the sign of electromagnetic chirality of time-harmonic polarised solutions of Maxwell’s equations generated by the corresponding $\operatorname{curl}$ eigenpair (see (E.6), (E.7)). We refer the reader to Appendix E for further details and a self-contained exposition of this fact. These are but few of the reasons why a more detailed examination of the spectrum of curl — and in particular of its spectral asymmetry — is very timely, if not overdue.
@@
and we denote by $|\,\cdot\,|$ the Euclidean norm of vectors. Throughout this paper we use Greek letters for tensor indices.
@@
    The operators $_{0}, P+P_{+}$ and $P_{-}$ are pseudodifferential of order zero.
@@
    $(P_{0})_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)=\|\xi\|^{-2}\,\xi_{\alpha}\,g^{\beta\gamma}(x)\,\xi_{\gamma}\,$, (1.8)
    $(P_{\pm})_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)=\frac{1}{2}\left[\delta_{\alpha}{}^{\beta}-[(P_{0})_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)\pm i\,\|\xi\|^{-1}\,E_{\alpha}{}^{\gamma\beta}(x)\,\xi_{\gamma}\right],$ (1.9)
@@
Remarkably, the pseudodifferential projections $P_{\pm}$ can be constructed explicitly. In Proposition 5.3 we will provide an algorithm for the calculation of their full symbols. The latter proposition and Theorem 1.3 give a detailed description of the structure of the projection operators $P_{0}$ and $P_{\pm}$, which is our first main result.
@@
$\{\text{positive eigenvalues}\}\,-\,\#\{\text{negative eigenvalues}\}=\operatorname{Tr}(P_{+}-P_{-}).$ (1.12)
@@
The operator $A$ is prima facie a pseudodifferential operator of order 0. However, taking the matrix trace brings about unexpected cancellations, so that the operator $A$ turns out, in fact, to be of order 3-3. This is our second main result.
@@
The asymmetry operator $A$ is a pseudodifferential operator of order 3-3 and its principal symbol reads
@@
A sufficient condition for a self-adjoint pseudodifferential operator to be of trace class and to have continuous integral kernel is for its order to be strictly less than d-d, where $d$ is the dimension of the manifold. See also [59, §12.1]. In the case at hand, the operator $A$ has order 3-3 and the dimension of the manifold is 3, so we are looking at a borderline situation. One would hope that with a mild regularisation one could cross the finish line and be able to define a notion of regularised trace.
@@
Pseudodifferential projections are very natural tools in the study of the spectral properties of systems. At first glance, one may be tempted to think they are unnecessarily complicated, especially for operators acting on trivial bundles, in that one could try and diagonalise the operator along the lines of [14] instead, thus reducing the problem at hand to the examination a collection of scalar operators. However, after examining the matter more carefully, diagonalisation turns out to be topologically obstructed in numerous examples of physical relevance. In particular, it was shown in [16, Section 3.3] that for the case of the operator $\operatorname{curl}$ one cannot even diagonalise the principal symbol $\operatorname{curl}_{\mathrm{prin}}$ (2.1) in the whole cotangent fibre at a given point of $M$, let alone the operator itself.
@@
There are, of course, many different approaches to the subject of spectral asymmetry. A notable one, for instance, relies on a local invariant known as Wodzicki residue [64], see also [51, 13] for a broader discussion. Since the literature on spectral asymmetry is vast, we decided to refrain from including a detailed literature review here, as it would inevitably be incomplete and would deserve a separate paper (if not a monograph).
@@
In Section 2 we put $\operatorname{curl}$ on a rigorous operator-theoretic footing, and state its basic properties. Although most of the results from this section are known at least in some form, detailed proofs, often hard to find in the literature, are provided in Appendix B in full and in a self-contained fashion, for the convenience of the reader and future reference. In doing so, we shall rely on the auxiliary operator $\operatorname{curl}_{E}$, called extended curl, which can be viewed as an elliptic “extension” of $\operatorname{curl}$.
@@
In Section 5 we study the pseudodifferential projections $_{0}, P+P_{+}$ and $P_{-}$, and provide an explicit algorithm for the construction of their full symbols. This section contains the proof of Theorem 1.3.
In Section 6 we define the asymmetry operator $A$ (subsection 6.1), prove that it is a pseudodifferential operator of order 3-3 (subsection 6.2), and compute its principal symbol $A_{\mathrm{prin}}$ (subsection 6.3). This section establishes Theorem 1.4.
@@
The paper is complemented by a list of notation given at the end of this section, and by seven appendices. In Appendix A we summarise our notation and conventions on exterior calculus. In Appendix B we provide proofs for the basic properties of the operator curl, summarised in Section 2. Appendices C and D are concerned with spectral theory on Berger 3-spheres: in Appendix C we recall the definition of a Berger sphere and list the eigenvalues of the Laplace–Beltrami operator, whereas in Appendix D we list the eigenvalues of $\operatorname{curl}$ on a Berger sphere as an illustration of spectral asymmetry, and compute the eta invariant by explicitly examining the eta function for curl. In Appendix E we examine the relation between the spectral problem for curl and solutions of Maxwell’s equations, and discuss the physical meaning of the sign of the eigenvalues of $\operatorname{curl}$. In Appendix F we derive series expansions for the parallel transport maps. Finally, in Appendix G we independently verify formula (1.13).
@@
$\|\,\cdot\,\|$ Riemannian norm (1.7)
$|\,\cdot\,|$ Euclidean norm
$\langle\,\cdot\,\rangle$ Japanese bracket (7.5)
AA Asymmetry operator, Definition 6.1
@@
$\mathfrak{a}(x,y)$ Integral kernel of the asymmetry operator AA
@@
$\operatorname{curl}_{E}$ Extended curl, Definition B.1
$\operatorname{curl}_{E,\mathrm{d}},\operatorname{curl}_{E,\delta},\operatorname{curl}_{E,\mathcal{H}}$ Orthogonal summands of $\operatorname{curl}_{E}$, Lemma B.4
@@
[…]
@@
$(\mu_{j},f_{j})$, $j=0,1,2,\ldots$ Eigensystem for $\Delta$
@@
$Q_{\mathrm{prin},s}$ Principal symbol of $Q$, a pseudodifferential operator of order s-s
@@
TMTM, $T^{*}M$ Tangent and cotangent bundle
@@
$Z$ Parallel transport map (4.6), see also Appendix F
@@ 2 The operator curl
In this section we put $\operatorname{curl}$ on a rigorous mathematical foundation, defining it within the framework of the theory of unbounded self-adjoint operators in Hilbert spaces. This would provide the reader with an account of the fundamental properties of $\operatorname{curl}$, in a self-contained fashion, postponing proofs until Appendix B.
@@
$\operatorname{curl}_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)=-i\,E_{\alpha}{}^{\beta\gamma}(x)\,\xi_{\gamma}\,$, (2.1)
where the tensor $E$ is defined in accordance with (1.10). Although $\xi$ is not a 1-form in its own right, at a formal level the RHS of (2.1) can be rewritten, concisely, as $i*\xi$. A straightforward calculation shows that
@@
The eigenvalues of $\operatorname{curl}_{\mathrm{prin}}$ are simple and read
@@
The basic properties of the operator (2.4) are summarised by the following theorem.
@@
    The spectrum of $\operatorname{curl}$ is discrete and accumulates to $\infty$ and -$\infty$.
@@
Let us point out, once again, that throughout this paper $\operatorname{curl}$ comes in two guises: as a differential expression (1.5) and as a self-adjoint operator (2.4). Both will be denoted by $\operatorname{curl}$”; which is which will be clear from the context.